magnitude of vector 2a+3b if a= i+j-2k and b=2i-j+k

Grade:11

2 Answers

vikas askiitian expert
13 years ago

2a+3b = 2(i + j - 2k) + 3(2 i- j + k)

          =5i - j - k

magnitude = (52 + 12+12)1/2 = (27)1/2 units

Aniket Patra
13 years ago

The vector sum comes out to be 8i-j-k.So magnitude is nearly 8.12.